# Cosmological Infrared Subtractions & Infrared-Safe Computables

### v1

@@ III The IR behaviour of cosmological integrals
Figure 2: Example of nestrohedra. It encodes the asymptotic structure of the three-site line graph. Starting with the top-dimensional simplex, whose facets are associated to tubings, it can be realised as a sequential truncation of such a simplex based on the underlying subgraph. The faces of the nestohedron are associated then to overlapping tubings. This construction is implemented to each of the terms in which the integrand can be triangulated (on the left and center). The full convergence region is given by the overlap between the nestohedra associated to the chosen triangulation (on the right)
@@ IV.1 Tree-level subtractions
$\mathfrak{W}^{{}^{\prime}\mbox{\tiny$(j_{1})$}}\>=\>\begin{pmatrix}\lambda^{{}^{\prime}\mbox{\tiny$(j_{1})$}}\\ -\mathfrak{e}_{\mbox{\tiny$j_{1}$}}\end{pmatrix},\qquad\mathfrak{W}^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{\mathfrak{g}}})$}}\>=\>\begin{pmatrix}\mbox{min}\{\lambda^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{\mathfrak{g}}})$}}\}\\ \mathfrak{e}_{j_{1}\ldots j_{n_{s}^{\mbox{\tiny$(\mathfrak{g})$}}}}\end{pmatrix}$ (13)
with the $\mbox{min}\{\lambda^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{\mathfrak{g}}})$}}\}$ being the minimum taken among all the nestohedra for the covectors with the same $\mathfrak{e}_{j_{1}\ldots j_{n_{s}^{\mbox{\tiny$(\mathfrak{g})$}}}}$.
@@
Power-law divergences – Power-law divergences are a little more intricate. There is a larger number of sectors which contribute to the divergent behaviour of the integral, concretely the ones identified by the adjacent covectors $\{\mathfrak{W}^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{\mathfrak{g}}})$}}\}$ such that $\lambda^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{\mathfrak{g}}})$}}\geq 0$. Recall that there is a hierarchy among the $\lambda$’s such that is larger the one associated to larger $n_{s}^{\mbox{\tiny$(\mathfrak{g})$}}$. This implies that it is enough to identify the smaller positive $\lambda$ or which of them vanish to know all the covectors that bound the sectors containing the divergences. As mentioned above, for each sector the divergences can be extracted by representing the integrand via a Mellin-Barnes representation – for more detail see Benincasa:2024lxe as well as the two-site tree example in the appendix. This implies that, depending on the degree of divergences along the relevant directions. For integer values of (at least some of the $\lambda$’s) multiple poles in the regulator can arise. The general logic follows what discussed in the above paragraph about the logarithmic divergences: one considers the integrals with trivial numerators obtained via a canonical form triangulation of the cosmological polytope, for each of these integrals perform the asymptotic analysis via the underlying nestohedron and identifies the sectors containing the divergences, and for each sector single out the divergences via a (multiple) Mellin-Barnes representation of the integrand. The precise structure indeed changes depending on the degree of divergence and the sector contributing.
@@ IV.2 Loop-level subtractions
Let us consider our integrals just over the loop momenta, with the $\{x_{s}\,s\in\mathcal{V}\}$ parametrising the actual external kinematics – this occurs, for example, for the conformally coupled scalars with conformal interactions. One can construct the associated nestohedron and consider the covectors compatible with the integration contour $\Gamma$. Then, the previous statement about the asymptotic structure translates in the statement that: in the ultraviolet the covector $\mathfrak{W}^{\mbox{\tiny$(1\ldots n_{e})$}}$ is the only divergent direction, and the sectors defined through it are the only one contributing to such a divergence; in the infrared instead the directions $\{\mathfrak{W}^{{}^{\prime}\mbox{\tiny$(j)$}}\}$ corresponding to the same loop ought to be taken separately.
@@
Some comments are now in order. First, note that each triangle diagram depends just on three edge weights, rather than the original four, but the integration is still over the four edge weights: the integration along this missing edge weight can be performed returning the integration measure and the integration contour of a triangle graph – see Benincasa:2024lxe . Secondly, this thick graph stays ultraviolet finite both in the site and edge weight integrations. Furthermore, as anticipated, were we to take the flat-space limit, we would not obtain the known flat-space infrared safe combination among box and triangle diagrams: the collapsing operation implies a shift by one of the Mellin parameter in the integration measure for the weights associated to the two sites that have been collapsed (maintaining the correct dimensionality). So this procedure returns a computable that is infrared finite at a generic point in the physical kinematic regione but develops infrared divergences due to soft and collinear limits in the edge weights in the flat-space limit.
@@ Appendix A Explicit examples
Figure 3: Nestohedron associated to the two-site tree graph. It shows divergences along the normal vectors $\{\mathfrak{W}^{\mbox{\tiny$(12)$}},\,\mathfrak{W}^{\mbox{\tiny$(j)$}},\,\mathfrak{W}^{{}^{\prime}\mbox{\tiny$(j)$}},\,j=1,2\}$ that identify its facets. The potential IR divergences are captured by the sectors containing $\mathfrak{W}^{\mbox{\tiny$(12)$}}$, $\mathfrak{W}^{\mbox{\tiny$(1)$}}$, $\mathfrak{W}^{\mbox{\tiny$(2)$}}$.

### v2

@@ III The IR behaviour of cosmological integrals
Figure 2: Example of nestrohedra. It encodes the asymptotic structure of the three-site line graph. Starting with the top-dimensional simplex, whose facets are associated to tubings, it can be realised as a sequential truncation of such a simplex based on the underlying subgraph. The faces of the nestohedron are associated then to overlapping tubings. This construction is implemented to each of the terms in which the integrand can be triangulated (on the left and center). The full convergence region is given by the overlap between the nestohedra associated to the chosen triangulation (on the right) Benincasa:2024lxe .
@@ IV.1 Tree-level subtractions
$\mathfrak{W}^{{}^{\prime}\mbox{\tiny$(j_{1})$}}\>=\>\begin{pmatrix}\lambda^{{}^{\prime}\mbox{\tiny$(j_{1})$}}\\ -\mathfrak{e}_{\mbox{\tiny$j_{1}$}}\end{pmatrix},\qquad\mathfrak{W}^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{(\mathfrak{g})}})$}}\>=\>\begin{pmatrix}\mbox{min}\{\lambda^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{\mathfrak{g}}})$}}\}\\ \mathfrak{e}_{j_{1}\ldots j_{n_{s}^{\mbox{\tiny$(\mathfrak{g})$}}}}\end{pmatrix}$ (13)
with the $\mbox{min}\{\lambda^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{(\mathfrak{g})}})$}}\}$ being the minimum taken among all the nestohedra for the covectors with the same $\mathfrak{e}_{j_{1}\ldots j_{n_{s}^{\mbox{\tiny$(\mathfrak{g})$}}}}$.
@@
Power-law divergences – Power-law divergences are a little more intricate. There is a larger number of sectors which contribute to the divergent behaviour of the integral, concretely the ones identified by the adjacent covectors $\{\mathfrak{W}^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{(\mathfrak{g})}})$}}\}$ such that $\lambda^{\mbox{\tiny$(j_{1}\ldots j_{n_{s}^{(\mathfrak{g})}})$}}\geq 0$. Recall that there is a hierarchy among the $\lambda$’s such that is larger the one associated to larger $n_{s}^{\mbox{\tiny$(\mathfrak{g})$}}$. This implies that it is enough to identify the smaller positive $\lambda$ or which of them vanish to know all the covectors that bound the sectors containing the divergences. As mentioned above, for each sector the divergences can be extracted by representing the integrand via a Mellin-Barnes representation – for more detail see Benincasa:2024lxe as well as the two-site tree example in the appendix. This implies that, depending on the degree of divergences along the relevant directions. For integer values of (at least some of the $\lambda$’s) multiple poles in the regulator can arise. The general logic follows what discussed in the above paragraph about the logarithmic divergences: one considers the integrals with trivial numerators obtained via a canonical form triangulation of the cosmological polytope, for each of these integrals perform the asymptotic analysis via the underlying nestohedron and identifies the sectors containing the divergences, and for each sector single out the divergences via a (multiple) Mellin-Barnes representation of the integrand. The precise structure indeed changes depending on the degree of divergence and the sector contributing.
@@ IV.2 Loop-level subtractions
Let us consider our integrals just over the loop momenta, with the $\{x_{s},\,s\in\mathcal{V}\}$ parametrising the actual external kinematics – this occurs, for example, for the conformally coupled scalars with conformal interactions. One can construct the associated nestohedron and consider the covectors compatible with the integration contour $\Gamma$. Then, the previous statement about the asymptotic structure translates in the statement that: in the ultraviolet the covector $\mathfrak{W}^{\mbox{\tiny$(1\ldots n_{e})$}}$ is the only divergent direction, and the sectors defined through it are the only one contributing to such a divergence; in the infrared instead the directions $\{\mathfrak{W}^{{}^{\prime}\mbox{\tiny$(j)$}}\}$ corresponding to the same loop ought to be taken separately.
@@
Some comments are now in order. First, note that each triangle diagram depends just on three edge weights, rather than the original four, but the integration is still over the four edge weights: the integration along this missing edge weight can be performed returning the integration measure and the integration contour of a triangle graph – see Benincasa:2024lxe . Secondly, this thick graph stays ultraviolet finite both in the site and edge weight integrations. Furthermore, as anticipated, were we to take the flat-space limit, we would not obtain the known flat-space infrared safe combination among box and triangle diagrams: the collapsing operation implies a shift by one of the Mellin parameter in the integration measure for the weights associated to the two sites that have been collapsed (maintaining the correct dimensionality). So this procedure returns a computable that is infrared finite at a generic point in the physical kinematic region but develops infrared divergences due to soft and collinear limits in the edge weights in the flat-space limit.
@@ Appendix A Explicit examples
Figure 3: Nestohedron associated to the two-site tree graph. It shows divergences along the normal vectors $\{\mathfrak{W}^{\mbox{\tiny$(12)$}},\,\mathfrak{W}^{\mbox{\tiny$(j)$}},\,\mathfrak{W}^{{}^{\prime}\mbox{\tiny$(j)$}},\,j=1,2\}$ that identify its facets. The potential IR divergences are captured by the sectors containing $\mathfrak{W}^{\mbox{\tiny$(12)$}}$, $\mathfrak{W}^{\mbox{\tiny$(1)$}}$, $\mathfrak{W}^{\mbox{\tiny$(2)$}}$ Benincasa:2024lxe .